• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 16:07:20 on July 28, 2006

    Entry number 175714

    keyword=Shift Summary (July 28 DAY)

    08:30 Stopped run 1763/20287
    08:30 FPP 1.5in C IN
    08:32 Started run 1764/20288 (still in ep elastic mode)
    09:07 Stopped run 1764/20288 for an access to fix the radiator control
    11:50 Hall is being locked up
    12:10 FPP 1.5in C OUT
    12:36 MCC called. Some problem with the optics. Expect (Mike Tieffenbach)
    has been called in
    12:53 Beam back. Target moving to LH2 position
    13:04 Run 1766/20289 started.
    13:13 MCC is going to pause the beam very briefly. Stopped runs.
    13:16 Beam back. Started runs 1767/20290 started. Set limit of number or
    counts at 2.5M for left DAQ to avoid file splitting.
    13:57 Stopped runs 1767/20290. MCC wants to check the beam stability for
    the weekend.
    14:12 MCC finished test. During the test, they were able to deliver up to
    20 uA without any problem. Started runs 1768/20291.
    14:43 Stopped run 1768. Limit on number of events reduced to 2M.
    14:43 Started 1769. Right DAQ still run 20291.
    15:07 Stopped run 1769 (automatic stop at 2M without file split)
    15:09 Started run 1770

    Summary of runs for ep elastic at kinematic point #2
    1763 06:57 08:23 86min
    1764 08:32 09:08 36min (122min)
    1766 13:05 13:15 10min (132min)
    1767 13:17 13:58 41min (173min)
    1768 14:13 14:42 29min (202min)
    1769 14:44 15:07 23min (225min)
    1770 15:09 15:37 28min (253min)

    15:37 Stopped run 1770 finishing ep elastic at 0% (4 hour run)
    15:37 S0 put IN
    15:38 Stopped R-DAQ run 20291
    15:39 Downloaded new trigger
    15:47 Changed prescale factor PS3=4 to have 500Hz of trigger 3. Trigger 6
    seems to be around 2kHz, so PS6 has been left to 1
    15:49 Started run 1771/20292
    16:00 At the end of shift, we are wondering huge negative deadtime with
    L-DAQ (something like -174%???)